The speed of a stream is 3mph. A boat travels 8 miles upstream in the same time it takes to travel 14 miles downstream. What is the speed of the boat in still water?
Answer by rajagopalan(174)

Let stream speed = 3 mph(given)
Let Boat speed in still water - x mph
Then
speed of boat up Stream=x-3
speed of boat down stream=x+3
then
Time taken for up stream journey=dist/speed=8/(x-3)
Time taken for down stream journey=dist/speed=14/(x+3)
But both these times are equal
equating 8/(x-3)=14/(x+3)
8(x+3)=14(x-3)
8x+24=14x-42
42+24=6x
66=6x
11=x
x=11
The speed of boat in still water =11 mph
And this verifies the condition upstream 8 miles in one hour and
down stream 14 miles in 14/(11+3)=14/14=1 hour
Ok